# Decoupled Classifier-Free Guidance for Counterfactual Diffusion Models

### v2

@@ D.1 Dataset details
For the CelebA-HQ dataset [20], we select Smiling, Male, and Young as three independent binary variables. These are among the most reliably annotated attributes in CelebA, each achieving over 95% consistency in manual labeling [60]. Moreover, they exhibit low inconsistency across duplicate face pairs (e.g., Male: 0.005; Smiling: 0.077), suggesting minimal label noise. We assume these variables to be independent, as our goal is to isolate and analyze attribute amplification under global classifier-free guidance (CFG), which is more interpretable with uncorrelated factors. As shown in Fig. A.1, the Pearson correlation matrix confirms weak pairwise correlations among these attributes. Although a moderate negative correlation is observed between Male and Young ($\rho=-0.33$), we attribute this to dataset bias rather than a true causal dependency, and proceed by modeling them as independent.
@@
D.2 Extra quantitative result for Celeba-HQ
Table A.3: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility deteriorates with increasing $\omega_{\mathrm{aff}}$, DCFG consistently maintains better reversibility than global CFG with $\omega=\omega_{\mathrm{aff}}$.
do(key) CFG Smiling AUC/$\Delta$ Male AUC/$\Delta$ Young AUC/$\Delta$ MAE LPIPS
do(Smiling) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
@@ D.2 Extra quantitative result for Celeba-HQ
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 89.9 / +3.4 95.8 / -1.1 77.8 / -0.8 0.128 0.093
@@
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 85.1 / -1.5 91.3 / -0.5 79.0 / -0.8 0.137 0.097
@@
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 87.4 / -0.1 95.0 / -0.7 63.2 / +0.9 0.129 0.095
@@
Table A.4: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Smiling AUC/$\Delta$ Male AUC/$\Delta$ Young AUC/$\Delta$ MAE LPIPS
do(Smiling) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 99.1 / +12.6 92.6 / -4.3 75.2 / -3.4 0.165 0.118
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 83.4 / -3.2 99.4 / +7.6 68.9 / -10.9 0.173 0.122
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 83.4 / -4.1 85.9 / -9.8 85.1 / +22.8 0.169 0.127
@@
D.3 Extra qualitative results for CelebA-HQ
Figure A.2: Additional qualitative results for do(Male) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves non-intervened attributes and identity while effectively reflecting the intervention.
@@ D.3 Extra qualitative results for CelebA-HQ
Figure A.3: Additional qualitative results for do(Smiling) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
@@
Figure A.4: Additional qualitative results for do(Young) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
@@
Figure A.5: Reversibility analysis for do(Similing) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
@@
Figure A.6: Reversibility analysis for do(Male) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
@@
Figure A.7: Reversibility analysis for do(Young) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. DCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of DCFG in enhancing both targeted editability and reversibility.
@@
Figure A.8: Effect of $\omega_{\text{inv}}$ on CelebA-HQ counterfactuals. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega=2.5$) and our proposed DCFG with fixed intervention guidance ($\omega_{\text{aff}}=2.5$) and varying invariant guidance $\omega_{\text{inv}}\in\{1.0,1.5,2.0,2.5\}$. As $\omega_{\text{inv}}$ increases, amplification of invariant attributes becomes more pronounced, and at $\omega_{\text{inv}}=2.5$, DCFG effectively reproduces the same over-editing behavior as global CFG. This shows that $\omega_{\text{inv}}$ modulates the degree of guidance applied to invariant attributes and should be carefully calibrated to maintain identity and disentanglement.
Appendix E EMEBD
E.1 Dataset details
@@
We use the EMory BrEast imaging Dataset (EMBED) [18] for our experiments. Schueppert et al. [48] manually labeled 22,012 images with circular markers and trained a classifier on this subset, which was then applied to the full dataset to infer circle annotations. We adopt this preprocessing pipeline and extract the circle attribute from their predictions. To define the density label, we binarize the original four-category breast density annotations by grouping categories A and B as low density, and categories C and D as high density. While the full dataset comprises 151,948 training, 7,156 validation, and 43,669 test samples, we use only 1,000 test samples in this work due to the high computational cost of diffusion models. As shown in Fig. A.9, the Pearson correlation matrix reveals that density and circle are nearly uncorrelated, supporting our assumption of their independence.
@@
E.2 Extra quantitative results for EMBED
Table A.5: EMBED: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{aff}}$. Compared to global CFG (i.e., $\omega$), DCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we apply $\omega_{\mathrm{inv}}=1.2$ to prevent degradation of invariant attributes. While reversibility deteriorates with increasing $\omega_{\mathrm{aff}}$, DCFG consistently maintains better reversibility than global CFG with $\omega=\omega_{\mathrm{aff}}$.
do(key) CFG Density AUC/$\Delta$ Circle AUC/$\Delta$ MAE LPIPS
do(density) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
@@ E.2 Extra quantitative results for EMBED
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 73.1 / +9.7 92.8 / -0.1 0.028 0.038
@@
$\omega_{\mathrm{aff}}=1.2$, $\omega_{\mathrm{inv}}=1.0$ 93.3 / +0.7 92.6 / +2.0 0.024 0.028
@@
Table A.6: EMBED: Effectiveness (ROC-AUC $\uparrow$) and Reversibility (MAE, LPIPS $\downarrow$) metrics when changing $\omega_{\text{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega=2.5$.
do(key) CFG Density AUC/$\Delta$ Circle AUC/$\Delta$ MAE LPIPS
do(density) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 96.7 / +33.3 89.5 / -3.4 0.035 0.063
@@
$\omega_{\mathrm{aff}}=2.5$, $\omega_{\mathrm{inv}}=1.0$ 92.2 / -0.4 98.5 / +7.9 0.028 0.039
@@
E.3 Extra qualitative results for EMBED
Figure A.10: Additional qualitative results for do(density) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in circle count, whereas DCFG mitigates such amplification.
Figure A.11: Additional qualitative results for do(circle) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and DCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$) and the difference map (CF-input). DCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in density, whereas DCFG mitigates such amplification.
Figure A.12: Reversibility analysis for do(density) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). DCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Figure A.13: Reversibility analysis for do(circle) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or DCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). DCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Appendix F MIMIC
F.1 Dataset details
We use the MIMIC-CXR dataset [19] in our experiments. Following the dataset splits and filtering protocols of [9, 13], we focus on the binary disease label for pleural effusion. We adopt the same causal graph (DAG) as proposed in De Sousa Ribeiro et al. [9], in which age is modeled as a parent of finding. While we include age as part of the conditioning variables, we do not intervene on it. Instead, our primary goal is to study amplification of unintervened variables caused by CFG. For this purpose, we focus on race, sex, and finding, which we assume to be mutually independent. Fig. A.14 shows the Pearson correlation matrix of these three attributes, where all pairwise correlations are small (e.g., $\rho=0.12$ between race and sex, and $\rho=-0.15$ between race and finding), supporting the validity of the independence assumption in our counterfactual modeling.

### v5

@@ F.1 Dataset details
For the CelebA-HQ dataset (Karras et al., 2017), we select Smiling, Male, and Young as three independent binary variables. These are among the most reliably annotated attributes in CelebA, each achieving over 95% consistency in manual labeling (Wu et al., 2023). Moreover, they exhibit low inconsistency across duplicate face pairs (e.g., Male: 0.005; Smiling: 0.077), suggesting minimal label noise. We assume these variables to be independent, as our goal is to isolate and analyze attribute amplification under global classifier-free guidance (CFG), which is more interpretable with uncorrelated factors. As shown in Fig. A.1, the Pearson correlation matrix confirms weak pairwise correlations among these attributes. Although a moderate negative correlation is observed between Male and Young ($\rho{=}-0.33$), we attribute this to dataset bias rather than a true causal dependency, and proceed by modeling them as independent.
@@
F.2 Extra quantitative result for Celeba-HQ
Table A.7: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{aff}}$. Compared to global CFG, FCFG achieves strong intervention effectiveness on intervened attributes while mitigating amplification on invariant attributes. For larger $\omega_{\mathrm{aff}}$, we set $\omega_{\mathrm{inv}}{=}1.2$ to prevent degradation of invariant attributes. Across all interventions, FCFG maintains better reversibility than global CFG with matched guidance strength.
[…]
do(Smiling) CFG (Ho and Salimans, 2022) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
@@ F.2 Extra quantitative result for Celeba-HQ
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 89.9 / +3.4 95.8 / -1.1 77.8 / -0.8 0.128 0.093
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 85.1 / -1.5 91.3 / -0.5 79.0 / -0.8 0.137 0.097
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 87.4 / -0.1 95.0 / -0.7 63.2 / +0.9 0.129 0.095
@@
Table A.8: CelebA-HQ: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}{=}2.5$, the amplification on invariant attributes becomes comparable to that of the global CFG setting with $\omega{=}2.5$.
[…]
do(Smiling) CFG (Ho and Salimans, 2022) $\omega=1.0$ 86.5 / +0.0 96.9 / +0.0 78.6 / +0.0 0.113 0.082
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 99.1 / +12.6 92.6 / -4.3 75.2 / -3.4 0.165 0.118
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 83.4 / -3.2 99.4 / +7.6 68.9 / -10.9 0.173 0.122
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 83.4 / -4.1 85.9 / -9.8 85.1 / +22.8 0.169 0.127
@@
F.3 Extra qualitative results for CelebA-HQ
@@
Figure A.2: Additional qualitative results for do(Smiling) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
@@
Figure A.3: Additional qualitative results for do(Male) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
Figure A.4: Additional qualitative results for do(Young) on CelebA-HQ. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while effectively reflecting the intervention.
@@
Figure A.5: Reversibility analysis for do(Similing) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed FCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. FCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of FCFG in enhancing both targeted editability and reversibility.
@@
Figure A.6: Reversibility analysis for do(Male) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed FCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. FCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of FCFG in enhancing both targeted editability and reversibility.
@@
Figure A.7: Reversibility analysis for do(Young) on CelebA-HQ. Each row shows the original image, followed by counterfactuals generated using global CFG ($\omega$) and our proposed FCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), along with their respective reversed generations. FCFG more faithfully preserves non-intervened attributes, resulting in visually and semantically more consistent reversals. This highlights the benefit of FCFG in enhancing both targeted editability and reversibility.
@@
Figure A.8: Effect of $\omega_{\text{inv}}$ on CelebA-HQ counterfactuals. Each row shows the original image followed by counterfactuals generated using global CFG ($\omega{=}2.5$) and our proposed FCFG with fixed intervention guidance ($\omega_{\text{aff}}{=}2.5$) and varying invariant guidance $\omega_{\text{inv}}\in\{1.0,1.5,2.0,2.5\}$. As $\omega_{\text{inv}}$ increases, amplification of invariant attributes becomes more pronounced, and at $\omega_{\text{inv}}{=}2.5$, FCFG effectively reproduces the same over-editing behavior as global CFG. This shows that $\omega_{\text{inv}}$ modulates the degree of guidance applied to invariant attributes and should be carefully calibrated to maintain identity and disentanglement.
Appendix G EMEBD
G.1 Dataset details
@@
We use the EMory BrEast imaging Dataset (EMBED) (Jeong et al., 2023) for our experiments. Schueppert et al. (2024) manually labeled 22,012 images with circular markers and trained a classifier on this subset, which was then applied to the full dataset to infer circle annotations. We adopt this preprocessing pipeline and extract the circle attribute from their predictions. To define the density label, we binarize the original four-category breast density annotations by grouping categories A and B as low density, and categories C and D as high density. While the full dataset comprises 151,948 training, 7,156 validation, and 43,669 test samples, we use only 1,000 test samples in this work due to the high computational cost of diffusion models. As shown in Fig. A.9, the Pearson correlation matrix reveals that density and circle are nearly uncorrelated, supporting our assumption of their independence.
@@
G.2 Extra quantitative results for EMBED
Table A.9: EMBED: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{aff}}$. Compared to global CFG (i.e., $\omega$), FCFG achieves strong intervention effectiveness on the intervened variable while mitigating amplification on invariant variables. For higher $\omega_{\mathrm{aff}}$, we set $\omega_{\mathrm{inv}}{=}1.2$ to prevent degradation of invariant attributes. While reversibility deteriorates with increasing $\omega_{\mathrm{aff}}$, FCFG consistently maintains better reversibility than global CFG with $\omega{=}\omega_{\mathrm{aff}}$.
[…]
do(density) CFG (Ho and Salimans, 2022) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
@@ G.2 Extra quantitative results for EMBED
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 73.1 / +9.7 92.8 / -0.1 0.028 0.038
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=1.2,\ \omega_{\mathrm{inv}}=1.0$ 93.3 / +0.7 92.6 / +2.0 0.024 0.028
@@
Table A.10: EMBED: Effectiveness (ROC-AUC $\uparrow$) and reversibility (MAE, LPIPS $\downarrow$) when varying $\omega_{\mathrm{inv}}$. Increasing $\omega_{\mathrm{inv}}$ consistently increases effectiveness on invariant variables, while degrading intervention effectiveness. When $\omega_{\mathrm{inv}}=2.5$, the amplification on invariant attributes becomes comparable to the global CFG setting with $\omega=2.5$.
[…]
do(density) CFG (Ho and Salimans, 2022) $\omega=1.0$ 63.4 / +0.0 92.9 / +0.0 0.027 0.033
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 96.7 / +33.3 89.5 / -3.4 0.035 0.063
@@
FCFG (Ours) $\omega_{\mathrm{aff}}=2.5,\ \omega_{\mathrm{inv}}=1.0$ 92.2 / -0.4 98.5 / +7.9 0.028 0.039
@@
G.3 Extra qualitative results for EMBED
Figure A.10: Additional qualitative results for do(density) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$). FCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in circle count, whereas FCFG mitigates such amplification.
Figure A.11: Additional qualitative results for do(circle) on EMBED. Each row shows the original image followed by counterfactuals generated with global CFG ($\omega$) and FCFG ($\omega_{\text{aff}},\omega_{\text{inv}}$) and the difference map (CF-input). FCFG better preserves invariant attributes and identity while effectively reflecting the intervention. Notably, under global CFG, increasing $\omega$ leads to spurious changes in density, whereas FCFG mitigates such amplification.
Figure A.12: Reversibility analysis for do(density) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or FCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). FCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Figure A.13: Reversibility analysis for do(circle) on EMBED. Each row shows the original image, the counterfactual generated using global CFG ($\omega$) or FCFG ($\omega_{\text{int}},\omega_{\text{inv}}$), their corresponding reversed generations, and the associated difference maps (counterfactual - input, and reversed - input). FCFG more faithfully preserves non-intervened attributes and leads to smaller residuals in the difference maps, indicating better identity preservation.
Appendix H MIMIC
H.1 Dataset details
We use the MIMIC-CXR dataset (Johnson et al., 2019) in our experiments. Following the dataset splits and filtering protocols of (Ribeiro et al., 2023; Glocker et al., 2023), we focus on the binary disease label for pleural effusion. We adopt the same causal graph (DAG) as proposed in Ribeiro et al. (2023), in which age is modeled as a parent of finding. While we include age as part of the conditioning variables, we do not intervene on it. Instead, our primary goal is to study amplification of unintervened variables caused by CFG. For this purpose, we focus on race, sex, and finding, which we assume to be mutually independent. Fig. A.14 shows the Pearson correlation matrix of these three attributes, where all pairwise correlations are small (e.g., $\rho{=}0.12$ between race and sex, and $\rho{=}-0.15$ between race and finding), supporting the validity of the independence assumption in our counterfactual modeling.